Please help me solve the following problem straight from a worksheet:
Use the law of exponents to find teh value of a taht would make the following stsatement true. Explain your reasoning.
|x+a| = b
4^16 = 16^2x = 256^x = 4,294,967,296

Answer by Alan3354(69443)

4^16 = 16^2x = 256^x = 4,294,967,296
--------------
4^16 = 16^2x
16 = 4^2, so 16^2x = 4^4x
=============
4^16 = 4^4x, so 4x = 16
x = 4
Substitute it, and it's correct.
--------------------
BTW, 4,294,967,296 = 2^32 which is 4 Gigabytes, or Gigahertz, Giga whatever.
2^10 = 1024 = 1 KBytes
2^20 = 1,048,576 = 1 Meg
2^30 = 1,073,741,824 = 1 Gig